OpenStudy (anonymous):

Goofy Golf ticket prices for kids under 12 years of age increased from $4.00 to $6.00 this year. What is the percent of increase in the tickets? 50% 60% 40%

OpenStudy (anonymous):

im not very good at percentages

OpenStudy (anonymous):

Price was :4 New price: 6 Amt of increase 6-4 = 2 % of increase 2/4 = 0.5 = 50% Hope this helps! :)

OpenStudy (anonymous):

6/4= 1.5*100 150% increase

OpenStudy (anonymous):

thank you ZhangYan